# Tunable Topological Phases in Quantum Kirigamis

### v1

@@ I Wave packet evolution and estimation of $r(\theta),t(\theta)$
Figure S5: $M$ dependence: $M$ dependence of $r(\theta)$ and $t(\theta)$ in BHZ junction modelling (see Fig.2(a) of main text) with ($N=1250$, $\alpha=1,\beta=1$ and $w=5$).
II Parameter dependence of $r(\theta)$ and $t(\theta)$
@@
System size ($N$) dependence: With the fixed set of parameters $(\alpha,\beta,w$ and $M$), we model the junction using both the microscopic BHZ junction and the effective $1d$ non-hermitian system. We find that the functional forms of $r(\theta)$ and $t(\theta)$ are independent of the system size ($N$) as shown in Fig. S2.
@@
Coupling dependence $(\alpha,w)$: Given a set of parameters ($N,M,w$, and $\beta$) we find that tuning the coupling strength ($\alpha$) can lead a quantitative change of $r(\theta)$ and $t(\theta)$, as illustrated in Fig. S3. The smaller value of $\alpha$ reduces the junction strength impeding transmission thus resulting in $r(\theta)>t(\theta)$. Interestingly, we note that for $\alpha=0.4$, the junction shows $r=t$ when $\theta\approx 0$ and $\frac{\pi}{2}$. This corresponds to the critical angles $(\theta_{1c}$ and $\theta_{2c})$ (as mentioned in main text). By tuning $\alpha$ to $1.2$ one can tune these critical angles up to $\theta\approx\frac{\pi}{4}$ (Fig. S3(c)) Further, increase in $\alpha$ makes $t(\theta)>r(\theta)$ as shown in Fig. S3(d). Similarly tuning $w$ can allow further modulation of $r(\theta)$ and $t(\theta)$ as shown in Fig. S4. Thus $\alpha,w$ serve as important tuning parameters in the system.
Figure S6: $M^{*}$ physics: (a) $r$ and $t$ as function of $M$ when $\theta=0$ with $N=2450$ and $N=4050$ with ($\alpha=1,\beta=1,w=5$). (b) $r$ and $t$ as function of $M$ when $\theta=\frac{\pi}{6}$ with $N=1800$ and $N=2450$ with ($\alpha=1,\beta=1,w=5$)
@@
$M$ dependence: The localization length $\xi$ of the edge states is related to $M$. Deep in a topological phase when $M=-3$ and a given $w=5$ the system has $\xi\ll w$ thus it allows easy transmission with $(t>r)$ as shown in Fig. S5(a) for all values of $\theta$. Note as $M$ is decreased further to $M=-3.6$ and $-3.8$, $\xi(M)$ increases since the bulk gap closing happens at $M=-4$. Given $t\sim 1$ near $\theta\sim 0$ implies that $\xi(M)<w$. However now in an intermediate range of $\theta$ near $\frac{\pi}{4}$ we find $r>t$. This increase in $r$ can be attributed to an effective reduction of junction coupling due to a coupling strength ($\lambda(\theta)$).
@@
System size dependence of $M^{*}$: In Fig. S6 we show the system size dependence of $r,t$ as a function of $M$ for a fixed $\theta=0,\frac{\pi}{6}$. We find that the critical point $M^{*}$ (where $r=t$) is independent of the system size. We further note that near $M\sim-4$, where the bulk gap closes, system size needs to be significantly large to reduce finite size effects.
Figure S7: Median lattice: (a) Unit cell of Kirigami structure. (b) Median lattice of a unit cell of Kirigami structure.
Figure S8: Edge state profile:  $P_{y}$ as function of $y$ when $L_{y}=16,r=0.4,t=0.916$. For (I) $k_{x}=-0.5,\epsilon=0.125$ and for (II) $k_{x}=0.5,\epsilon=0.125$
III $U_{\text{eff}}$ and edge states
In order to calculate the effective time evolution operator $U_{\text{eff}}$ of the complete network (see eqn. (4) of main text) one needs to form a median lattice of the Kirigami network. Such a lattice is constructed by replacing every side of the Kirigami block with a site placed in the middle and with hopping amplitudes that are unidirectional and follow the same convention of $r(\theta)$ and $t(\theta)$ as in the scattering junctions. The unit cell of the original Kirigami network and the corresponding median lattice is shown in Fig. S7. Fourier transforming this non-Hermitian Hamiltonian leads to the eigenvalues $e^{-i\epsilon}$, from which $\epsilon$ can be obtained. A similar calculation for the triangular motif leads to a six-site unit cell. The same problems are solved on a ribbon geometry, with periodic boundary conditions in the $x$ direction to give us the band dispersions as reported in the main text. Our treatment closely follows Charles et al. (2020).
Now we investigate the edge state profile of diagonalizing $U_{\text{eff}}$ on a ribbon geometry which has translational symmetry in the $x$ direction and is finite in the $y$ direction. The spectrum is shown in Fig. 3(d) and (f) in the main text. The electronic probability density $P_{y}$ (as a function of $y$) for given values of $k_{x}$ and $\epsilon$ is shown in Fig. S8. As can be seen, the right (left) moving edge state resides on the bottom (top) of the ribbon. These states are absent when periodic boundary conditions are used in the $y$ direction.

### v2

@@ IX Appendix
Figure 8: System size ($N$) dependence:  The plot of $r(\theta)$ and $t(\theta)$ in BHZ junction modelling (see Fig.3) when $N=450$ and $N=1250$ with $\alpha=1,\beta=1,w=5$ and $M=-3.6$.
Figure 9: Coupling dependence:  $r(\theta)$ and $t(\theta)$ in BHZ junction modelling (see Fig.3) with ($N=1250$, $M=-3.6,\beta=1$ and $w=5$). (a) $\alpha=0.4$ corresponds to the $r=t$ when $\theta\approx 0$ or $\frac{\pi}{2}$ and $r\approx 1$, $t\approx 0$ when $\theta\approx\frac{\pi}{2}$. (b) For $\alpha=1.2$, $r=t$ when $\theta\approx\frac{\pi}{4}$. (c) For $\alpha=1.5$, $t(\theta)>r(\theta)$.
Figure 10: $w$ dependence: The junction width $(w)$ dependence of $r(\theta)$ and $t(\theta)$ in BHZ junction modelling (see Fig.3) with ($N=1250$, $\alpha=1,\beta=1$ and $M=-3.6$).
Figure 11: $M$ dependence: $M$ dependence of $r(\theta)$ and $t(\theta)$ in BHZ junction modelling (see Fig.3) with ($N=1250$, $\alpha=1,\beta=1$ and $w=5$).
Appendix A Parameter dependence of $r(\theta)$ and $t(\theta)$
@@
System size ($N$) dependence: With the fixed set of parameters $(\alpha,\beta,w$ and $M$), we model the junction using both the microscopic BHZ junction modelling (eqn. (4)) and the ENHJ modelling (eqn. (13)). With the wavepacket evolution, we find that the functional forms of $r(\theta)$ and $t(\theta)$ are independent of the system size ($N$) as shown in Fig. 8 (not shown for ENJH model).
@@
Coupling dependence $(\alpha,w)$: Given a set of parameters ($N,M,w$, and $\beta$) we find that tuning the coupling strength ($\alpha$) can lead a quantitative change of $r(\theta)$ and $t(\theta)$, as illustrated in Fig. 9. The smaller value of $\alpha$ reduces the junction strength impeding transmission thus resulting in $r(\theta)>t(\theta)$. Interestingly, we note that for $\alpha=0.4$, the junction shows $r=t$ when $\theta\approx 0$ and $\frac{\pi}{2}$ (Fig. 9(a)). This corresponds to the critical angles $(\theta_{1c}$ and $\theta_{2c})$. By tuning $\alpha$ to $1.2$ one can tune these critical angles up to $\theta\approx\frac{\pi}{4}$ (Fig. 9(b). Further, increase in $\alpha$ makes $t(\theta)>r(\theta)$ as shown in Fig. 9(c). Similarly tuning $w$ can allow further modulation of $r(\theta)$ and $t(\theta)$ as shown in Fig. 10. Thus $\alpha,w$ serve as important tuning parameters in the system.
@@
$M$ dependence: The localization length $\xi$ of the edge states is related to $M$. Deep in a topological phase when $M=-3$ and a given $w=5$ the system has $\xi\ll w$ thus it allows easy transmission with $(t>r)$ as shown in Fig. 11(a) for all values of $\theta$. Note as $M$ is decreased further to $M=-3.6$, $\xi(M)$ increases since the bulk gap closing happens at $M=-4$. Given $t\sim 1$ near $\theta\sim 0$ implies that $\xi(M)<w$. However now in an intermediate range of $\theta$ near $\frac{\pi}{4}$ we find $r>t$. This increase in $r$ can be attributed to an effective reduction of junction coupling due to a coupling strength ($\lambda(\theta)$).
Appendix B Stretched Kirigami and non-uniform strain
Figure 12: Stretched Kirigami:  The time evolution of wavepacket for three configurations of the Kirigami networks at (a) $\theta_{o}=15^{\circ},\Delta_{\theta}=5^{\circ}$ for three different times $\tau=\{2,80,130\}$ (shown by arrows). (b) $\theta_{o}=40^{\circ},\Delta_{\theta}=15^{\circ}$ and $\tau=\{2,60,80\}$. (c) $\theta_{o}=85^{\circ},\Delta_{\theta}=13^{\circ}$ and $\tau=\{2,80,190\}$.
In order to study the effects on realistic strain, as seen in experiments and simulations [70], we model a strain gradient and see its effect on the transport properties. We assume translational invariance in the vertical ($y$) direction and assume that a node $n$ at any position $x_{n}\in\{-\frac{l^{\prime}}{2},\frac{l^{\prime}}{2}\}$ where $l^{\prime}$ is the size of the complete Kirigami network at an angle $\theta_{o}$ (see eqn. (2)). To include strain we further add a Gaussian profile such that any node has an angle $\theta_{n}=\theta_{o}-\delta\theta_{n}$ where
$\delta\theta_{n}=\frac{\Delta_{\theta}}{(e^{-1/2}-1)}\left[\exp\left(-\frac{2x_{n}^{2}}{l^{\prime 2}}\right)-1\right]$ (17)
This ensures that as we move from the center to the edge of the Kirigami networks the node angles smoothly deform from $\theta_{o}$ to $\theta_{o}-\Delta_{\theta}$ as seen in experiments on auxetic Kirigamis . Given the characterization of the node (see Fig. 3) a local angle of $\theta_{n}$ corresponds to a scattering coefficient $\{r_{n},t_{n}\}$. Thus $\{r_{n},t_{n}\}$ will vary from $\{r(\theta_{o}),t(\theta_{o})\}$ at the center to $\{r(\theta_{o}-\Delta_{\theta}),t(\theta_{o}-\Delta_{\theta})\}$ at the edges of the Kirigami network. We denote the maximum deviation in $r_{n}$ from center to edge as $\Delta_{r}=r(\theta_{o})-r(\theta_{o}-\Delta_{\theta}).$ Assuming $\Delta_{\theta}\ll\theta_{o}$ we linearly extrapolate the variation of $r_{n}$ to be $r_{n}=r(\theta_{o})-\delta r_{n}$ where
$\delta r_{n}=\frac{\Delta_{r}}{(e^{-1/2}-1)}\left[\exp\left(-\frac{2x_{n}^{2}}{l^{\prime 2}}\right)-1\right]$ (18)
such that at every node $|t_{n}|=\sqrt{1-r^{2}_{n}}$.
In Fig. 4 we had shown the character of the network when the node angles were uniform i.e. $\theta_{o}=15^{\circ},40^{\circ},85^{\circ}$. We now add a $\Delta_{\theta}=5^{\circ},15^{\circ},13^{\circ}$ in each of the cases and show the wavepacket evolution in Fig. 12. This shows that the results of the uniform strain case is stable to small gradients in the strain profile. This is however only true deep in the topological/trivial phase such that local $r_{n}$ at any node do not cross the phase boundaries. Strongly non-uniform strains ($\Delta_{\theta}\gg\theta_{o}$) may bring interesting transport properties which is an interesting future work.
Appendix C Correspondence between ENHJ and CC networks
In order to further corroborate the results between the ENHJ and the CC networks we compare the spectrum of the ENHJ and the CC network in a strip geometry with different values of $r$ in Fig. 13. For a given choice of $r(\theta),t(\theta)$ we find the eigenvalues of the evolution operator constructed using ENHJ formulation ($U_{\text{NH}}$) which are the form $e^{-i\chi}$ and plot $\chi$ as a function of $r$. With open boundary conditions, we find that close to $\chi=0$ and $r<r_{c}=\frac{1}{\sqrt{2}}$ the plot has “edge states” as shown in Fig. 13(a). These edge states disappear under the periodic boundary conditions as shown in Fig. 13(b). This clearly shows the presence of edge states and the topological character in the network. When $r>r_{c}=\frac{1}{\sqrt{2}}$, there are no edge states also, $\chi$ has a finite gap, which shows the trivial insulating phase of the network. Clearly, ENHJ formulation is also able to capture the topological phase transition and presence of critical point $r_{c}=\frac{1}{\sqrt{2}}$ in the CC network as discussed in section V.3.
Figure 13: Variation of $\chi$ with $r$ Plot of eigenvalues $\chi$ of $i\ln(U_{\text{NH}})$ as a function of $r$ with (a) open boundary conditions, there are edge states for $r<r_{c}=\frac{1}{\sqrt{2}}$. (b) With the periodic boundary conditions in both directions, all the edge states disappear for $r<r_{c}=\frac{1}{\sqrt{2}}$. The red dotted line corresponds with the critical point $r_{c}=\frac{1}{\sqrt{2}}$ obtained from CC formulation (see section V.3).
Appendix D System size dependence of $M^{*}$
Figure 14: $M^{*}$ physics: (a) $r$ and $t$ as function of $M$ when $\theta=0$ with $N=2450$ and $N=4050$ with ($\alpha=1,\beta=1,w=5$). (b) $r$ and $t$ as function of $M$ when $\theta=\frac{\pi}{6}$ with $N=1800$ and $N=2450$ with ($\alpha=1,\beta=1,w=5$).
@@ Appendix D System size dependence of $M^{*}$
In Fig. 14 we show the system size dependence of $r,t$ as a function of $M$ for a fixed $\theta=0,\frac{\pi}{6}$. We find that the critical point $M^{*}$ (where $r=t$) is independent of the system size. We further note that near $M\sim-4$, where the bulk gap closes, system size needs to be significantly large to reduce finite size effects.